OpenStudy (anonymous):

Solve -2x + 1 less than or equal to -7 and festive the graph of solution. A. X less than or equal to 4; closed circle on 4, shading to left. B. x greater than or equal to 4; closed circle on 4, shading to right. C. X greater than or equal to 4; open circle on 4, shading to right. D. X less than or equal to 4; open circle on 4, shading to left.

OpenStudy (anonymous):

Describe not festive*

OpenStudy (anonymous):

basically what you should be doing on this problem is solving for x. -2x+1=-7 solve for x. Then that should bascially give you you're answer :)

OpenStudy (anonymous):

But not sure if it's an open or closed circle..

OpenStudy (anonymous):

If the inequality includes the 'or equals to' qualifier, then it is always closed at some point. -2x + 2 <= -7 This solution will have a close circle.

OpenStudy (anonymous):

- 2x + 1 <= -7 -2x <= -7 -1 -2x <= - 8 -2x/-2 => -8/-2 x => 4 Answer : B...x is greater then or equal to four, closed circle on 4 shaded to the right :)

OpenStudy (anonymous):

Well done!